I need help on this expression.
16x-8=14x-2
The answer is 3, but I don't know how.

Answer by ikleyn(53751)

.
I need help on this expression.
16x-8=14x-2
~~~~~~~~~~~~~~~~~~~

16x - 8 = 14x - 2

Add -14x to both sides. You will get

16x - 14x - 8 = -2,  or

2x - 8 = -2.

Add 8 to both sides. You will get

2x = -2 + 8,  or

2x = 6.

Divide both sides by 2. You will get

x = 3.